OpenStudy (anonymous):

Hey I need help with some questions they should be pretty easy. Any and all help appreciated.

OpenStudy (anonymous):

\[\sqrt{-32}-\sqrt{-18}\] Find the exact value.

OpenStudy (anonymous):

Can you guys help me?

OpenStudy (anonymous):

@trueseminole @jcpd910

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

Okay, I think the exact value would be the square root of 32 i. Like this: 5.65i

OpenStudy (saifoo.khan):

Why the decimals? :c

OpenStudy (anonymous):

Because the square root of a negative number is imaginary. I've given up trying to understand why, it just is.

OpenStudy (anonymous):

Well he wants the exact value, so might as well try and get close.

OpenStudy (anonymous):

Alright But my prof says i need to show work :/

OpenStudy (saifoo.khan):

\[\sqrt{-32} = 4i \sqrt2 \]This is exact.

OpenStudy (anonymous):

:D thanks, how about the rest of the problem?

OpenStudy (saifoo.khan):

@jcpd910 : Ring a bell?

OpenStudy (anonymous):

Yes, but can you explain it to me?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (saifoo.khan):

\[\sqrt{-32} = i \sqrt{32} =i \sqrt{16*2}=i \sqrt{4*4*2}=4i \sqrt{2}\]

OpenStudy (anonymous):

Oh yeah, because sqrt -32 is sqrt -1 * 32!

OpenStudy (anonymous):

\[\sqrt{-32}-\sqrt{-18}=4 i \sqrt{2}-3 i \sqrt{2}=\sqrt{2} (4 i-3 i)=i \sqrt{2} \]

OpenStudy (saifoo.khan):

Yes, sir.

OpenStudy (anonymous):

But one thing, how does i sqrt 4*4*2 become 4i sqrt 2

OpenStudy (anonymous):

alright, so if i were to do the same thing for the rest of the equation

OpenStudy (saifoo.khan):

Whenever there's a pair of numbers inside the root, one of them comes out. Like \[\sqrt4 = \sqrt{2*2} = 2\]or \[\sqrt18 = \sqrt{3*3*2} = 3 \sqrt2 \]

OpenStudy (anonymous):

You beat me to it

OpenStudy (anonymous):

Okay, so can you throw me a practice problem?

OpenStudy (anonymous):

So would the \[\sqrt{-18}\] equal \[3i \sqrt{2}\] ?

OpenStudy (jdoe0001):

\(\bf \sqrt{-32}-\sqrt{-18} \\ \quad \\ {\color{brown}{ 32\to 2\cdot 2\cdot 2\cdot 2\cdot 2\to 2^4\cdot 2\to (2^2)^2\cdot 2}} \\ \quad \\ {\color{brown}{ 18\to 2\cdot 3\cdot 3\to 2\cdot 3^2}}\qquad thus \\ \quad \\ \sqrt{-32}-\sqrt{-18}\implies \sqrt{-1\cdot 32}-\sqrt{-1\cdot 18}\implies \sqrt{-1}\cdot \sqrt{32}-\sqrt{-1}\cdot \sqrt{18} \\ \quad \\ i\ \sqrt{(2^2)^2\cdot 2}-i\ \sqrt{2\cdot 3^2}\)

OpenStudy (anonymous):

@saifoo.khan hey just one more problem

OpenStudy (saifoo.khan):

Shoot.

OpenStudy (anonymous):

\[\frac{ 3 + 4i }{ 4- 2i}\] Simplify and Rationalize.

OpenStudy (saifoo.khan):

@jcpd910 wanna try?

OpenStudy (saifoo.khan):

\[\frac{3+4i}{4-2i} \times \frac{4+2i}{4+2i} \]

OpenStudy (anonymous):

Can you divide complex numbers? Like: 4i/-2i = -2i?

OpenStudy (anonymous):

Okay I'll try.

OpenStudy (anonymous):

And nevermind saifoo o_o But alright, once I have what you just did, how do I proceed?

OpenStudy (anonymous):

\[\frac{ (3+4i)(4+2i) }{ (4-2i)(4+2i) } \times \frac{ (3+4i)(4+2i) }{ (4-2i)(4+2i) }\] Right so far?

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

\[\frac{ (3+4i)(4+2i) }{ (4-2i)(4+2i) } \times \frac{ (3+4i)(4-2i) }{ (4-2i)(4+2i) }\]

OpenStudy (anonymous):

There, correct so far?

OpenStudy (anonymous):

How did you... I don't understand

OpenStudy (anonymous):

I multiplied both things by the other denominator in order to get common denominatros.

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

But if you do that wouldn't you only get the first half of what you got?

OpenStudy (saifoo.khan):

O_O

OpenStudy (anonymous):

Lol saifoo whats going on

OpenStudy (anonymous):

lol yeah derp

OpenStudy (anonymous):

I DONT EVEN NKOW!

OpenStudy (anonymous):

halp meeeee *slowly dies*

OpenStudy (anonymous):

@saifoo.khan Help us wise one.

OpenStudy (anonymous):

wait isn't is just: \[\frac{3+4i}{4-2i} \times1\]

OpenStudy (anonymous):

In which case isn't it just: \[\frac{3+4i}{4-2i} \]

OpenStudy (anonymous):

You'd end up with the same exact equation you started with though....Wouldnt you?

OpenStudy (anonymous):

Lol what?!

OpenStudy (saifoo.khan):

Lol. Okay so: \[\frac{ 3 + 4i }{ 4- 2i}\]Rationalize. \[\frac{3+4i}{4-2i} \times \frac{4+2i}{4+2i}\] Now multiply. \[\frac{(3+4i)(4+2i)}{(4-2i)(4+2i)}\]

OpenStudy (anonymous):

So you do nothing?!

OpenStudy (anonymous):

Ahhhh

OpenStudy (anonymous):

Oh I thought we started off different

OpenStudy (saifoo.khan):

Now just use FOIL.

OpenStudy (anonymous):

Lol jcpd you did some crazy **** I was wondering where you got so many numbers

OpenStudy (anonymous):

lerl brb

OpenStudy (anonymous):

And alright saifoo I'll try it.

OpenStudy (anonymous):

Alright does it look something like this:

OpenStudy (anonymous):